# DriveScape: Towards High-Resolution Controllable Multi-View Driving Video Generation

### v3

@@ Abstract
Recent advancements in generative models have provided promising solutions for synthesizing realistic driving videos, which are crucial for training autonomous driving perception models. However, existing approaches often struggle with multi-view video generation due to the challenges of integrating 3D information while maintaining spatial-temporal consistency and effectively learning from a unified model. In this paper, we propose an end-to-end framework named DriveScape for multi-view, 3D condition-guided video generation. DriveScape not only streamlines the process by integrating camera data to ensure comprehensive spatial-temporal coverage, but also introduces a Bi-Directional Modulated Transformer module to effectively align 3D road structural information. As a result, our approach enables precise control over video generation, significantly enhancing realism and providing a robust solution for generating multi-view driving videos. Our framework achieves state-of-the-art results on the nuScenes dataset, demonstrating impressive generative quality metrics with an FID score of 8.34 and an FVD score of 76.39, as well as superior performance across various perception tasks. This paves the way for more accurate environmental simulations in autonomous driving. Code will be available at our project homepage.

### v4

@@ Abstract
Recent advancements in generative models have provided promising solutions for synthesizing realistic driving videos, which are crucial for training autonomous driving perception models. However, existing approaches often struggle with multi-view video generation due to the challenges of integrating 3D information while maintaining spatial-temporal consistency and effectively learning from a unified model. We propose DriveScape, an end-to-end framework for multi-view, 3D condition-guided video generation, capable of producing 1024 x 576 high-resolution videos at 10Hz. Unlike other methods limited to 2Hz due to the 3D box annotation frame rate, DriveScape overcomes this with its ability to operate under sparse conditions. Our Bi-Directional Modulated Transformer (BiMot) ensures precise alignment of 3D structural information, maintaining spatial-temporal consistency. DriveScape excels in video generation performance, achieving state-of-the-art results on the nuScenes dataset with an FID score of 8.34 and an FVD score of 76.39. Code will be available at our project homepage.